OpenStudy (sneakysquash):

One file clerk can file 10 folders per minute. A second file clerk can file 11 folders per minute. How many minutes would the two clerks together take to file 672 folders?

OpenStudy (sneakysquash):

______minutes.

OpenStudy (sneakysquash):

Algebra 2.

OpenStudy (sneakysquash):

@SkylyerBailey @MissSmartiez @math_man21

OpenStudy (math_man21):

im in calculus 1 bud

OpenStudy (math_man21):

remind me something here

OpenStudy (error1603):

I think It will take 32 minutes. In 1 minute, they file 21 folders combined. 672/21 equals 32.

OpenStudy (math_man21):

wait 10+11/672

OpenStudy (math_man21):

then divide by how many ppl are there

OpenStudy (phi):

rate * time = amount (of work) if they don't slow each other down, you can add their rates 10 folders/ min + 11 folders /min = 21 folders / min 21 * t= 672 t= 672/21 = 32 minutes